suppose a put option with a $40 strike is also available with premium of $2.80. calculate your percentage return for the six-month holding period if the stock price declines to $36 per share
Answer #1

SIX MONTH HOLDING PERIOD RETURN IS ASKED.

IN PUT OPTION, PAY OFF WILL BE POSITIVE, IF PRICE ON EXPIRY(36) IS LOWER THAN STRIKE PRICE(40)

HERE, PAY OFF = 40-36 = 4

HE HAS PURCHASED PUT OPTION, SO HE HAS PAID PREMIUM OF 2.80.

SO NET PROFIT = PAY OFF - PREMIUM PAID = 4 -2.8 =1.2

SIX MONTH HOLDING PERIOD RETURN = NET PROFIT/INVESTMENT = 1.2/2.8 = 0.42857 = 42.86%

HERE INVESTMENT IS PREMIUM PAID.

ANSWER : 42.86% (THUMBS UP PLEASE)